please help me with solve this equation what is the solution set 8x-7=19x+15
Answer by Jk22(389)

8x-7=19x+15 | sub 8x on both side
-7=19x-8x+15
-7=11x+15 | sub 15 on both side
-7-15=11x
-22=11x | divide by 11 on both side
-2=x

Solution : x=-2

Verification : 8*(-2)-7=-16-7=-23 (lhs)
19*(-2)+15=-38+15=-(38-15)=-23 (rhs=lsh)